OpenStudy (babyslapmafro):

Please help me factor/simplify the following. ((x^3)+(3x^2)-12x+4)/((x^3)-4t)

OpenStudy (babyslapmafro):

\[(x ^{3}+3^{2}-12x+4)/(x ^{3}-4x)\]

OpenStudy (goten77):

give me a sec to type this out

OpenStudy (babyslapmafro):

No rush, thank you for the help

OpenStudy (goten77):

\[\frac{ x ^{3}+3x^{2}-12x +4 }{ x ^{3} -4x} = \] i assume thats what you ment to type out....... i just did it then noticed the 3^2 and i missed it up and gotta redo now lol

OpenStudy (babyslapmafro):

ahh yes, sorry for the mix-up...apologies

OpenStudy (goten77):

its k

OpenStudy (goten77):

wow srry i cant do this 4 some reason.... i keep missing up and im not posting a wrong answer

OpenStudy (goten77):

all i can really do tbh since im obviously doing something wrong is \[\frac{ x(x ^{2}+3x-12)+4 }{ x(x ^{2} -4)}\]

OpenStudy (goten77):

i for some reason forgot what to do lol srry

OpenStudy (anonymous):

\[\frac{ x^ 3 +3x^ 2 −12x+4 }{ x ^3 −4x } =\] First, did you notice that numerator have 4 terms, recall, if we have 4 terms what should we do?

OpenStudy (babyslapmafro):

I wish i knew

OpenStudy (goten77):

veramath show us how to do this... i kinda forgot XD

OpenStudy (anonymous):

Let's factor the numerator: \[x^3+3x^2-12x+4\] group fisrt two terms and group last to terms we have \[x^3+3x^2-12x+4=(x^3+3x^2)+(-12x+4)\] Factor our GCF from each group: \[x^3+3x^2-12x+4=(x^3+3x^2)+(-12x+4)=x^2(x+3)-4(x+3)=(x+3)(x^2-4)\] So \[x^3+3x^2-12x+4=(x+3)(x+2)(x-2)\] Factor the denominator: \[x^3-4x=x(x^2-4)=x(x+2)(x-2)\] I think now you can do the rest right?